Calculate the time required to deposit 0.37g of Silver from a silver Nitrate solution using a current of 2.5A ( Charge of Ag in is AgNO 3 is +1)


1
Expert's answer
2020-12-28T04:41:11-0500

If Farade constant = 96485 Cmol-1

Atomic mass of Ag = 107.5 gmol-1


  • Deposition occurs as,

Ag+(aq) + e ------> Ag(s) ---- (1)


  • No of Ag moles deposited ( nAg ) = "\dfrac{0.37g}{107.5gmol^{-1}}"

= "3.44\times10^{-3} mol"


  • As per the equation (1),

"n_e = n_{Ag}\\"

  • Using,

"Q =It,\\3.44\times10^{-3}mol\;\times \;96485\; Cmol^{-1}\;=\;2.5A\;\times\;t\\
t\;=\;\bf132.76s"